OpenStudy (anonymous):

Verify the identity. Cot[theta-(pi/2)=-tan(theta)

OpenStudy (jdoe0001):

\(\bf cos\left(\theta-\frac{\pi}{2}\right)=-tan(\theta) \\ \quad \\ \quad \\ cos\left(\theta-\frac{\pi}{2}\right)\implies cos(\theta)cos\left(\frac{\pi}{2}\right)+sin(\theta)sin\left(\frac{\pi}{2}\right) \\ \quad \\ \textit{what is the }sin\left(\frac{\pi}{2}\right)\quad ? \\ \quad \\ \textit{what is the }cos\left(\frac{\pi}{2}\right)\quad ?\)

OpenStudy (jdoe0001):

hmmm well. it helps if I used cotangent identity, rather than cosine =) hehe lemme rewrite it some

OpenStudy (jdoe0001):

\(\bf cot\left(\theta-\frac{\pi}{2}\right)=-tan(\theta) \\ \quad \\ \quad \\ cot\left(\theta-\frac{\pi}{2}\right)\implies \cfrac{1}{tan\left(\theta-\frac{\pi}{2}\right)}\implies \cfrac{1}{\frac{tan(\theta)-tan\left(\frac{\pi}{2}\right)}{1+tan(\theta)tan\left(\frac{\pi}{2}\right)}} \implies \cfrac{1} { \frac{\frac{sin(\theta)}{cos(\theta)}-\frac{sin\left(\frac{\pi}{2}\right)}{cos\left(\frac{\pi}{2}\right)}}{1+\frac{sin(\theta)}{cos(\theta)}\frac{sin\left(\frac{\pi}{2}\right)}{cos\left(\frac{\pi}{2}\right)}} } \\ \quad \\ \textit{what is the }sin\left(\frac{\pi}{2}\right)\quad ? \\ \quad \\ \textit{what is the }cos\left(\frac{\pi}{2}\right)\quad ?\)

OpenStudy (anonymous):

Uuuuuuhhhhhhh.......

OpenStudy (jdoe0001):

lemme recheck myself

OpenStudy (anonymous):

It might not be you, I'm just not good at this algebra/trig stuff.

OpenStudy (jdoe0001):

http://hyperphysics.phy-astr.gsu.edu/hbase/imgmth/trid1.gif <--- is all we're using, the tangent identity

OpenStudy (anonymous):

Okay, I figured that much, but I need help doing most of the steps in general. I just have no idea where to go from here.

OpenStudy (jdoe0001):

hmmm actually... that didn't yield much ....

OpenStudy (jdoe0001):

lemme check something

OpenStudy (jdoe0001):

gimme a few secs

OpenStudy (anonymous):

Okay

OpenStudy (jdoe0001):

\(\bf cot\left(\theta-\frac{\pi}{2}\right)=-tan(\theta) \\ \quad \\ \quad \\ cot\left(\theta-\frac{\pi}{2}\right)\implies \cfrac{cos\left(\theta-\frac{\pi}{2}\right)}{sin\left(\theta-\frac{\pi}{2}\right)} \implies \cfrac{cos(\theta)cos\left(\frac{\pi}{2}\right)+sin(\theta)sin\left(\frac{\pi}{2}\right)}{sin(\theta)cos\left(\frac{\pi}{2}\right)-cos(\theta)sin\left(\frac{\pi}{2}\right)} \\ \quad \\ \textit{what is the }sin\left(\frac{\pi}{2}\right)\quad ? \\ \quad \\ \textit{what is the }cos\left(\frac{\pi}{2}\right)\quad ?\)

OpenStudy (jdoe0001):

you should be able to get those from the Unit Circle

OpenStudy (anonymous):

Okay, uuummm. Still not sure. I did it on my calculator and it said sin(pi/2)=1 and cos(pi/2)=0. I'm at a double disadvantage because I'm naturally bad at math, and I don't think these concepts were explained well in my text book.

OpenStudy (jdoe0001):

ahemm well do you happen to have a Unit Circle? many online if you need one

OpenStudy (anonymous):

Yeah, there's a unit circle.

OpenStudy (jdoe0001):

but yes, the calculator is correct check your Unit Circle so we know that \(\bf sin\left(\frac{\pi}{2}\right)\implies 1\qquad cos\left(\frac{\pi}{2}\right)\implies 0\qquad thus \\ \quad \\ \cfrac{cos(\theta){\color{blue}{ cos\left(\frac{\pi}{2}\right)}}+sin(\theta){\color{blue}{ sin\left(\frac{\pi}{2}\right)}}}{sin(\theta){\color{blue}{ cos\left(\frac{\pi}{2}\right)}}-cos(\theta){\color{blue}{ sin\left(\frac{\pi}{2}\right)}}}\implies \cfrac{cos(\theta){\color{blue}{0}}+sin(\theta){\color{blue}{1}}}{sin(\theta){\color{blue}{ 0}}-cos(\theta){\color{blue}{1}}}\) so what does that leave you with?

OpenStudy (anonymous):

0+1=1. 0-1=-1. 1/-1=-1 ?

OpenStudy (jdoe0001):

? whatever happened to \(\bf sin(\theta)\ and \ cos(\theta)\quad ?\)

OpenStudy (anonymous):

Uuuuuhhhhh, they went on holiday? Um I really don't know. I get the sense that I should know this, but I'm kinda clueless. Sorry for my ignorance. I'm looking at a unit circle online, I'm hoping it will be more informative than my text book.

OpenStudy (jdoe0001):

hehe... lemme rewrite that

OpenStudy (jdoe0001):

thought surely they'd be happy to go on holiday =)

OpenStudy (jdoe0001):

\(\bf sin\left(\frac{\pi}{2}\right)\implies 1\qquad cos\left(\frac{\pi}{2}\right)\implies 0\qquad thus \\ \quad \\ \cfrac{cos(\theta)\cdot {\color{blue}{ cos\left(\frac{\pi}{2}\right)}}+sin(\theta)\cdot {\color{blue}{ sin\left(\frac{\pi}{2}\right)}}}{sin(\theta)\cdot {\color{blue}{ cos\left(\frac{\pi}{2}\right)}}-cos(\theta)\cdot {\color{blue}{ sin\left(\frac{\pi}{2}\right)}}}\implies \cfrac{cos(\theta)\cdot {\color{blue}{0}}+sin(\theta)\cdot {\color{blue}{1}}}{sin(\theta)\cdot {\color{blue}{ 0}}-cos(\theta)\cdot {\color{blue}{1}}}\) so... what would you be left with?

OpenStudy (anonymous):

My jacked-up-math-rejecting-brain is telling me that answer is -[sin(theta)/cos(theta)]

OpenStudy (jdoe0001):

yeap \((\bf \cfrac{cos(\theta)\cdot {\color{blue}{0}}+sin(\theta)\cdot {\color{blue}{1}}}{sin(\theta)\cdot {\color{blue}{ 0}}-cos(\theta)\cdot {\color{blue}{1}}}\implies \cfrac{sin(\theta)}{-cos(\theta)}\implies -\cfrac{sin(\theta)}{cos(\theta)}\) does that look like ... .an identity ?

OpenStudy (jdoe0001):

\(\bf -\cfrac{sin(\theta)}{cos(\theta)}\implies -1\cdot {\color{blue}{ \cfrac{sin(\theta)}{cos(\theta)}}}\)

OpenStudy (anonymous):

Okay, so is that what I'm looking for? -1 * sine(theta)/cos(theta)?

OpenStudy (jdoe0001):

http://www.efunda.com/math/trig_functions/images/trig_relation.gif <--- any of those look like it ?

OpenStudy (anonymous):

Yeah, first one at the top. tan=sinx/cosx

OpenStudy (jdoe0001):

ahhh well, there

OpenStudy (jdoe0001):

\(\bf \bf -\cfrac{sin(\theta)}{cos(\theta)}\implies -1\cdot {\color{blue}{ \cfrac{sin(\theta)}{cos(\theta)}}}\implies -{\color{blue}{ tan(\theta)}} \)

OpenStudy (anonymous):

Oh, cool. I feel like I've been tricked into getting the correct answer, but I'll take it. Thanks. But, stay available for a few minutes, will ya, I've got try and name the steps that get me from one step to the next, they're kinda strict about showing your work.

OpenStudy (jdoe0001):

well... sure... you can just check above, is pretty straightforward

OpenStudy (anonymous):

Cool, thanks.